Find the slope of the curve and the equation of tangent line of the parametric equation to the given point.



x=ln t , y = t ^ - 1 , when t = 2

1
Expert's answer
2021-11-29T16:25:13-0500

Given that "x=ln(t),y=t^{-1}"

"\therefore\frac{dx}{dt}=\frac{1}{t}, \frac{dy}{dt}=-\frac{1}{t^2}"

So, "\frac{dy}{dx}=\frac{\frac{dy}{dt}}{\frac{dx}{dt}}\\"

"\Rightarrow \frac{dy}{dx}=-\frac{1}{t}"

At "t=2, \frac{dy}{dx}=-\frac{1}{2}"

The slope of curve at "t=2" is "-\frac{1}{2}".

At "t=2, x= ln(2),y=\frac{1}{2}"

We can write:

"t=e^x,t=\frac{1}{y}"

Therefore, equation of curve is: "y=e^{-x}"

Equation of tangent at "t=2" is:

"y-\frac{1}{2}=\frac{1}{2}(x-ln(2))"

"\therefore y=\frac{x}{2}+\frac{1-ln(2)}{2}"